### v1

# CasTex: Cascaded Text-to-Texture Synthesis via Explicit Texture Maps and Physically-Based Shading

Abstract
This work investigates text-to-texture synthesis using diffusion models to generate physically-based texture maps. We aim to achieve realistic model appearances under varying lighting conditions. A prominent solution for the task is score distillation sampling. It allows recovering a complex texture using gradient guidance given a differentiable rasterization and shading pipeline. However, in practice, the aforementioned solution in conjunction with the widespread latent diffusion models produces severe visual artifacts and requires additional regularization such as implicit texture parameterization. As a more direct alternative, we propose an approach using cascaded diffusion models for texture synthesis (CasTex). In our setup, score distillation sampling yields high-quality textures out-of-the box. In particular, we were able to omit implicit texture parameterization in favor of an explicit parameterization to improve the procedure. In the experiments, we show that our approach significantly outperforms state-of-the-art optimization-based solutions on public texture synthesis benchmarks.
Project page: https://thecrazymage.github.io/CasTex/.
1 Introduction
The creation of 3D assets is a crucial part of modern computer graphics applications, with demand being supported by such industries as movie production, industrial design, and computer gaming, to name a few. High-quality assets should simultaneously meet aesthetic requirements and technical production requirements (i.e., high-quality mesh topology and limited polygon count). As a result, designing a 3D model is a labor-intensive creative task that involves multiple skills and is hard to automate due to the lack of high-quality data and occasionally difficulty in formalizing target criteria.
At the same time, the rapid development of generative models in the visual domain has already tremendously impacted various creative domains and even extends to 3D graphics. The successes in 3D reconstruction [25, 1, 17] and foundation text-to-image model [33, 32] led fully-automated end-to-end 3d asset creation prototypes [29, 23, 20]. Such solutions use the prior knowledge acquired by the diffusion model to generate 3D assets meeting aesthetic requirements. However, the resulting assets come in a volumetric representation and are often incompatible with modern graphics engines. Even with conversion tools for extracting meshes from volumetric representation, the assets end up having overly detailed geometry with occasional high-level geometric inconsistencies (e.g., multiple faces, as in the Janus effect). The aforementioned limitations may be critical when the technical requirements come into play.
Texture synthesis is a more modest attempt to automate 3D asset creation, aiming to create a realistic texture for a given 3D shape. The task assumes that all of the technical requirements regarding the model geometry are already addressed by an artist. To meet modern industry standards, in our work, we specifically consider the synthesis of physically based textures [4]. Such textures present a versatile instrument able to capture a diversity of materials and create realistically looking assets within varying lighting scenarios.
Some early texture synthesis approaches involved the back-projection of synthetic images on 3D shapes. However, such approaches are incompatible with PBR textures as they only allow the capture of a fraction of material properties through RGB colors. As an alternative, we consider an optimization based approach that allows capturing the material properties with back-propagation through the rendering engine. In particular, we continue the line of work studying various applications of Score Distillation Sampling [29].
We aimed to design a zero-shot texture-synthesis solution without relying on auxiliary 3D data and foundation model fine-tuning. However, we note that the texture synthesis setup unveils the limitations of Score Distillation Sampling applied in conjunction with modern diffusion models. Our observations go in line with the results of [41]. We continue their analysis and attribute the effect to the specifics of latent diffusion models.
As an alternative, we consider cascaded diffusion models that guide the images in RGB space rather than latent space. Our solution follows the pipeline of cascaded models. First, we generate a coarse texture using SDS guidance with a low-resolution model. Then we refine the texture using SDS guidance with a super-resolution model. Below, we summarize our main contributions.
  • We study the limitations of Score Distillation Sampling when applied to latent diffusion models. Our analysis indicates that the optimal solutions of the algorithm include images outside of the natural image domain.
  • Motivated by our analysis, we introduce a texture synthesis approach using Score Distillation Sampling with cascaded diffusion models. The proposed approach mimicks the diffusion model architecture and acts in two stages. The first stage generates a coarse texture, while the second stage improves overall quality of the texture obtained at the first stage.
  • We show that our approach outperforms state-of-the-art optimization based texture synthesis solutions on the Objaverse dataset. Score Distillation Sampling in pixel space allows omitting additional regularizers required in previous works. Additionally, we report the scaling performance of our method with the diffusion model size, the impact of lighting conditions used during training, and the role of physically based textures.
2 Related Work
Generating high-quality textures for 3D objects is a long-standing challenge in computer graphics and machine learning. Early approaches are based on category-specific generative models, such as generative adversarial networks (GANs) like Texture Fields [28], SPSG [11], LTG [42], and Texturify [36]. These methods often struggle with limited generalization and required extensive manual adjustments. Recent advances in large-scale text-to-image diffusion models [32, 33] transform texture synthesis, allowing more flexible and realistic results. However, different approaches vary in how they leverage diffusion models to incorporate 3D information and ensure global texture consistency.
Diffusion-based texture generation. A significant part of recent works formulates texture generation as an iterative inpainting process using depth-conditioned diffusion models. Text2Tex [6], TEXTure [31], and Paint3D [43] generate textures by synthesizing single-view images and progressively refining the texture using different object views. Although these zero-shot approaches allow for training-free generation, they often struggle with seams at visibility boundaries and suffer from baked-in lightning artifacts, leading to global inconsistencies. To mitigate these issues, some methods introduce additional refinement steps: Text2Tex employs an inpainting diffusion model for post-processing, while Paint3D refines textures directly in UV space. TexDreamer [22] follows a different approach by directly generating a UV texture map using a fine-tuned diffusion model, avoiding the need for iterative view synthesis and improving consistency across the surface.
An alternative strategy to address inconsistency is to introduce additional structures to enforce coherence between views. SyncMVD [21] tackles this by generating each view independently while employing a synchronization module that aligns outputs in every denoising step of the diffusion model. Similarly, TexFusion [5] suggest a module that aggregates multi-view object predictions into a unified latent texture representation, iteratively refining it throughout the diffusion process to ensure consistency across the entire object.
Another approach, inspired by MVDream [34], leverages consistent multi-view image generation as an initialization step for texture synthesis. Meta 3D TextureGen [2] follows this idea by generating four consistent views, which are projected onto a 3D surface to form an initial texture for subsequent refinement in UV space. MVPaint [8] extends this pipeline by introducing a spatial-aware 3D inpainting stage between projection and UV refinement, further enhancing coherence across different object views.
Optimization-based methods. Optimization-based techniques refine a texture by optimizing an objective function that evaluates rendered views. Early approaches such as CLIP-Mesh [26] and Text2Mesh [24] use differentiable rendering and CLIP [30] as a guidance signal to iteratively update texture parameters. However, these methods often produce blurry textures due to the lack of high-frequency details.
More recent works incorporate diffusion models to improve texture optimization. Latent-Paint [23], Fantasia3D [7], and Paint-it [41] combine differentiable rendering with the Score Distillation Sampling (SDS) technique [29], distilling knowledge from a diffusion model to refine textures. While this approach enhances realism, it often suffers from color oversaturation and high-frequency artifacts. FlashTex[13] mitigates some of these issues by employing a fine-tuned multi-view model to generate an initial textured view of the object, followed by an optimization stage using L2-loss and SDS for refinement. In this work, we continue the line of work employing SDS. However, we attribute some of the artifacts to the underlying latent diffusion model and propose using a pixel-space model. Aiming at physically-plausible textures, we use a text-to-image model pretrained on natural images and resort from using synthetic data for multi-view model adaptation.
PBR texture generation. Most of the works discussed above generate only RGB, yielding limited photo-realism. The works of Fantasia3D, Paint-it, and FlashTex propose to synthesize more advanced physically based textures [4], which are able to mimic different materials and look realistic under different lighting. We asses a similar approach to generation of PBR textures with a different generative model.
3 Preliminaries
In our work, we rely on a pretrained text-to-image model to generate textures. To apply a model designed for image generation to a different task, we use Score Distillation Sampling (SDS). Next, we describe the algorithm and the rendering algorithm required to apply it.
3.1 Score Distillation Sampling for Gradient Based Generation
SDS gained popularity in text-to-3D setup. In [29], the authors consider a differentiable rendering mechanism that produces a random view of a 3D model. In particular, they parameterize the model with a volumetric radiance field and render it using random lighting and camera position. SDS leverages a pretrained text-to-image diffusion model to guide the model parameters with gradient descent. Crucially, the mild assumptions on the rendering algorithm allow one to apply the method to other differentiable image generation mechanisms. In later works, Score-Distillation Sampling was successfully applied to other rendering functions and image parameterizations such as Gaussian Splatting [40, 9, 37], vector graphics [39, 14], and image rasterization with physically-based shading [20, 7, 41].
SDS defines gradients used to update the parameter $\theta\in\Theta$. Assume a mapping $x_{0}=g(\theta,z)$ that takes the parameter and a random variable $c$ and produces an image $x_{0}$. Given a diffusion process with noisy images $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ at time $t$ and scale-parameter $\bar{\alpha}_{t}$, the approach leverages a pretrained diffusion model $\epsilon_{\phi}(x_{t};y,t)$ to define a pseudo-objective used to update $x_{0}$. [^1] Throughout the paper, we assume that the diffusion model $\epsilon_{\phi}$ incorporates the condition $y$ through classifier-free guidance. The pseudo-objective $\hat{x}_{0}^{t}=\tfrac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\phi}(x_{t},t,y)}{\sqrt{\bar{\alpha}_{t}}}$ estimates the appearance of the rendered image $x_{0}$ based on a noisy sample $x_{t}$ with a single step diffusion model prediction.
The resulting gradient is estimated as follows:
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}(\theta)=\mathbb{E}_{t,\epsilon,c}\left[w(t)(x_{0}-\hat{x}_{0}^{t})\frac{\partial x_{0}(\theta,c)}{\partial\theta}\right],$ (1)
where $w(t)$ is an adjustable scalar weight parameter. Note that [29] formulate the method in noise prediction $\epsilon$-terms, while we introduce the equivalent formulation in $x_{0}$-terms. On the one hand, the gradient is essentially the mean-squared loss gradient $\mathbb{E}\left[w(t)\|x_{0}-\hat{x}_{0}^{t}\|^{2}\right]$ with the omitted dependence of $\hat{x}_{0}^{t}$ on $\theta$. On the other hand, the analysis in [29] demonstrates that the procedure minimizes $\mathcal{L}_{SDS}(\theta)=\mathbb{E}_{t,c}\operatorname{KL}\left(q(x_{t}\mid t)\|p_{\phi}(x_{t}\mid y,t)\right)$ using the score $\epsilon_{\phi}(x_{t};y,t)=\nabla_{x_{t}}\log p_{\phi}(x_{t}\mid y,t)$.
Originally, SDS was applied to a pixel-space diffusion model that generates low-resolution images. However, pretrained pixel-space models are less common, and, as a result, applications of the method were primarily focused on latent diffusion models (LDMs). These models approximate the distribution of the low-dimensional embeddings of a high-resolution images rather than the images itself. Crucially, the mapping $\text{enc}(\cdot)$ that produces the embedding is differentiable, as a result, can be considered as a component of a function generating an image:
$g_{\textit{latent}}(\theta,c)=\text{enc}(g_{\textit{pixel}}(\theta,c)).$ (2)
The potential advantage of latent diffusion models is the ability to guide high-resolution images compared to pixel-space models. At the same time, it requires additional resources to run forward and backward passes of the encoder network. In Section 4, we analyze other implications of applying SDS in latent space rather than pixel space.
3.2 Differentiable Physically Based Shading
Our texture synthesis approach applies SDS in a setting, where $g(\theta,z)$ renders an image of a fixed 3D model given a set of texture parameters $\theta$ and a random camera position and scene lighting specified by $z$. For rendering, we follow the pipeline proposed by [27].
In particular, we take a differentiable rasterization algorithm based on the Nvdiffrast library by [19]. For textures, we adopt the Disney-physically based material model from [4], allowing us to incorporate the obtained textures into popular graphics engines without further modifications. The texture $\theta=(\theta_{d},\theta_{r},\theta_{m},\theta_{n})$ consists of a diffuse component $\theta_{d}\in\mathbb{R}^{h\times w\times 3}$, a roughness map $\theta_{r}$ parameterizing a specular lobe and a metalness factor $\theta_{m}$ used to interpolate between plastic and metallic appearance. Finally, a normal map $\theta_{n}$ stores surface normal information, allowing to simulate additional geometric detail even within low-polygon models.
To model lighting we incorporate cube environment map with a split sum approximation for the outgoing radiance integral approximation [16].
Figure 1: Overall pipeline. Our method consists of two stages. On the first stage, given a 3D mesh and a prompt our method employs a differentiable rendering pipeline to generate random views of the model under various lighting. We update the randomly initialized model texture using Score Distillation Sampling (SDS). On the second stage, refine the texture from the first stage using SDS with a super-resolution diffusion model. For the same view and lighting, we render two frames: a frame with the fixed texture taken from the first step and a frame with a current texture. Using the former frame as the condition, we back-propagate the SDS gradients through the latter frame.
Figure 2: Optimization based image generation for a toy single-mode diffusion model. We study the effect of image parameterization and the difference between pixel and latent space diffusion models. Score Distillation Sampling (SDS) finds the distribution mode for both latent and pixel space diffusion models. Top row: the pixel space model achieves near-perfect reconstruction quality measured by PSNR. The latent code $\text{enc}(\theta_{\text{RGB}})$ is also close to the mode and leads to high quality reconstruction. However, for the latent model the optimized parameter $\theta_{\text{RGB}}$ fails to approximate the original image. Bottom row: The implicit regularization proposed in [41] significantly improves the results for SDS for the latent diffusion model, but still falls short of the image reconstruction given by the decoder or pixel space diffusion model.
4 Methodology
In the following section, we start with a motivating example and then describe the proposed method. The overall scheme is illustrated in Figure 1.
4.1 Limitations of Score Distillation Sampling for Latent Diffusion Models
In this section, for latent diffusion models, we argue that the procedure is prone to producing samples with artifacts due to the ill-posed nature of the task.
Following the common practice, we convey the analysis on a task of image generation. Specifically, we use a mapping $g_{id}(\theta,c)=\theta$ that outputs a 2D image $\theta$ and does not rely on the second stochastic variable. The goal is to generate images on par with the diffusion model samples with a gradient descent algorithm. Intuitively, the SDS gradients guide images $g(\theta,c)$ towards the manifold of natural images. Using the chain rule, these gradients also propagate to the parameter $\theta$. However, in the case of latent diffusion models guidance happens within the latent space: $g(\theta,c)=\text{enc}(g_{id}(\theta,c))$. As a result, the training signal must propagate through the encoder.
Next, to mitigate some of the potential issues of SDS in our analysis, we consider a toy diffusion model that represents a degenerate distribution $\delta(x-x^{*})$ concentrated at $x^{*}$. The problems may include a poorly trained score estimation model or the excessive variance of the stochastic gradient 1. For the degenerate distribution, we use the optimal model that separates the Gaussian noise from the original sample
$\epsilon_{degenerate}(x_{t})=\tfrac{x_{t}-\sqrt{\bar{\alpha}_{t}}x^{*}}{\sqrt{1-\bar{\alpha}_{t}}}.$ (3)
Note that for the model from Eq. 3 and constant weights $w(t)=1$ the gradient in Eq. 1 simplifies to the deterministic gradient of the mean squared error:
$\nabla_{\theta}\mathcal{L}_{SDS}(\theta)=\nabla_{\theta}\|g(\theta)-x^{*}\|^{2}.$ (4)
In other words, in the specified toy setup, we can assume
$\mathcal{L}_{SDS}(\theta)=\|g(\theta)-x^{*}\|^{2}$ (5)
To analyze SDS, we ran the optimization in various setups and report the results in Figure 2. For the pixel-space diffusion model with $g(\theta)=\theta$, gradient descent rapidly converges to $\theta\approx x^{*}$ (top row, left). For the latent-space model we assume $x^{*}$ to be the latent code of the same image and minimize $\|\text{enc}(g(\theta))-x^{*}\|^{2}$. Although gradient descent converges and the decoder maps $\text{enc}(\theta)$ to the original image (top row, right), the optimized parameter $\theta$ itself is a poor approximation to the original image (top row, middle). In practice, SDS with latent diffusion models finds globally optimal solutions that fall outside of the natural image domain.
Notably, the above observation does not limit text-to-3D applications. However, in the above toy setup and texture synthesis domain the observed artifacts limit the applicability of SDS. To avoid the artifacts, a recent work [41] proposed to parameterize textures with a deep convolutional network. Their solution builds upon the observations of [38] and acts as an implicit regularizer. It helps avoiding solutions outside of the natural image domain, but requires running an additional convolutional and, as we show below, still leads to noticeable image artifacts.
In the bottom row of Figure 2, we report SDS performance with an implicit image parameterization from [41]. We denote the parameterization as DIP (Deep Image Prior), following [38]. In this case, the image is an output of a UNet network with weights $\theta_{dip}$ and a fixed input sampled from the standard Gaussian distribution. In the context of the SDS, we can interpret the forward pass of the UNet as the differentiable rendering algorithm $g_{\text{UNet}}(\theta_{dip})$ constrained to generate natural images.
We minimize Eq. 5 for the implicit parameterization assuming both latent and pixel-space models. The implicit parameterization significantly improves the result for the latent diffusion model (middle row). The artifacts present in the explicit parameterization are less noticeable. Still, even with the implicit parameterization SDS in the latent space is unable to match the reconstruction quality of the reconstructed latent code (middle and right bottom row images). In our preliminary experiments, we tried to impose additional regularization on $\theta$ using the decoder outputs, but did not achieve noticeable improvements compared to the default SDS. Besides that, pixel-space optimization achieves significantly better result for both explicit and implicit image parameterization.
Given the above observations, in our work, we use SDS with pixel-space models.
4.2 Applying SDS with Super-Resolution Diffusion Models
The downside of pixel-space diffusion models is that the model for high-resolution images require significant resources. While latent diffusion models employ a VAE [18] to encode a high-resolution image into a low-dimensional latent code, the typical solution for pixel-space models is to generate an image in a cascaded fashion. At the first stage, diffusion model generates a low-resolution image. During the second stage, a super-resolution model generates a high-resolution version of the image from the first stage.
In [29], the authors run SDS only with a low-resolution model. However, the low-resolution model may be unable to recover fine-grained details and high-frequency features as its training domain was limited to down-scaled images. To achieve highly detailed textures, in our work we mimic the architecture of cascaded diffusion models. First, we run SDS with a low-resolution diffusion model. Then, we refine the output of the first stage with SDS combined with a super-resolution model.
Figure 3: Score distillation sampling applied to a super-resolution diffusion model. Horizontal axis represents progress over time. Top row: the same image $g(\theta)$ used both as a diffusion model input $x$ and the condition $x_{cond}$ for super-resolution. The resulting procedure is unstable and significantly diverges from the original image. Bottom row: same procedure with fixed super-resolution condition $x_{cond}=g(\theta_{0})$. The condition anchors the dynamics to the initial image.
In text-to-image setup, the second stage model takes a low-resolution image as an input along with a text prompt and time. This leads to the two natural ways of applying SDS to refine the first stage output. The first option is to pass the rendered image $g(\theta,c)$ as the diffusion model argument in SDS and a down-scaled copy of the image as a condition. Intuitively, in this case, the algorithm gradually updates the texture by adding the details lacking after the previous gradient step. The second option is to freeze the parameter $\theta_{0}=\theta$ obtained at the first stage and run SDS using a copy of $\theta$ as the new parameter and the previous $\theta_{0}$ as a condition. In this case, the first stage output in the condition acts as an anchor for the gradient descent dynamics.
In our experience, the procedure in the first solution tends to be less stable and leads toward out-of-domain images, whereas the second approach tends to improve the resolution while preserving the details of the first stage result. We illustrate our observations in Figure 3.
In summary, we propose a texture synthesis approach based on a cascaded modification of the SDS. Figure 1 is a schematic representation of the overall approach.
5 Experiments
We evaluated our method on the Objaverse dataset [12], comparing it both qualitatively and quantitatively with state-of-the-art texture synthesis approaches. We also performed an ablation study to validate key design decisions.
5.1 Evaluation setup
For our method, we took DeepFloyd-IF [35] as an open-source cascaded diffusion model. We considered the XL ($64{\times}64$) and L ($256{\times}256$) models for the first and second stages, respectively. Also, we considered the setting with the smaller model IF(M) at the first stage to explore the effect of the diffusion model scale on the texture quality.
We compared the proposed method with three baselines that represent various approaches to texture synthesis. First, Text2Tex [6] is a well-established back-projection-based method generating non-PBR textures. Second, Paint-it [41] is a simplistic optimization-based method using SDS. The crucial distinction between our method and Paint-it is the underlying diffusion model and texture parameterization. Finally, FlashTex [13] is another recent approach using SDS. It addresses the limitations of previous methods by fine-tuning the model used for guidance on multi-view 3D data with controlled lighting. Again, all of the baselines are based on latent diffusion models for texture synthesis.
To evaluate the quality of generated textures, we computed the Frechet Inception Distance (FID) [15] and Kernel Inception Distance (KID) [3] metrics on a subset of the Objaverse dataset [12]. Specifically, we adopted the same curated subset used in Text2Tex [6], which includes 410 high-quality textured meshes from 225 distinct categories. This subset was originally filtered to exclude meshes with simplistic or inconsistent textures, category mismatches, and over-triangulated or scanned objects. All methods generated textures for the same set of untextured 3D meshes using identical text prompts, then each textured mesh is rendered from 20 fixed viewpoints at a resolution of $512\times 512$ with a white background using Blender [10]. We provide a detailed evaluation setup in Appendix A, where we specify the most critical rendering parameters (e.g., lighting configuration, render engine, etc.) to ensure accurate comparisons.
5.2 Main Results
Figure 4: Qualitative comparison. We synthesized textures for models from the Objaverse dataset using the proposed approach and a number of recent competing methods. Our method generates seamless textures with softer colors compared with latent diffusion-based approaches.
Figure 5: Human preference study comparing our method with competing optimization based and back-projection baselines.
Size FID KID Runtime
($\downarrow$) ($\downarrow$) (×$\times 10^{-3}$, $\downarrow$) (seconds, $\downarrow$)
Text2Tex 2.5B 21.3 4.1 549
Paint-it 1B 25.2 6.7 892
FlashTex 1.3B 29.0 9.5 280
CasTex (M) 400M 24.9 5.6 45
CasTex (M+L) 1.6B 23.2 5.0 628
CasTex (XL) 4.3B 22.8 4.4 174
CasTex (XL+L) 5.5B 20.5 3.1 710
Table 1: Comparison of FID and KID scores for different text-to-texture generation methods. Notably, even with the smallest diffusion model, our method outperforms optimization-based baselines. As expected, the super-resolution stage improves the results for both stages. Surprisingly, in our evaluation, the older back-projection-based method outperforms the optimization-based method and is only rivaled by our largest setup. We use NVIDIA A100 80GB for time measurements and generate textures with a resolution of 1024x1024 pixels.
Automated quality assessment. We report FID and KID scores in Table 1. Our smallest single-stage setup outperformed the two optimization-based approaches. The two-stage cascaded diffusion approach achieved the best performance, outperforming all baselines. We illustrate the two stages in Figure 6 and provide additional illustrations in Appendix C. The second stage significantly refines textures, enhancing fine-grained details and coherence.
Notably, the weak performance of FlashTex on the benchmark is the result of the limited generalization abilities of the Light ControlNet proposed by the authors. To our surprise, Text2Tex delivered competitive results. Even though the method is the oldest baseline, the back-projected photo-realistic images lead to photo-realistic textures and renders.
Our method also achieves competitive synthesis times, with the two-stage variant completing generation in 12 minutes per mesh – comparable to baselines while delivering better quality. For time-constraint applications, the single-stage variant generates plausible textures in just 3 minutes, offering a compelling trade-off between computational efficiency and texture realism, making it suitable for time-constraint applications.
User preference study. In addition, we conduct a human study to compare the perceptual quality of our generated textures against baseline methods. The evaluation is conducted by professional assessors, asked to make a decision between two rendered textures given a textual prompt. The decision is made according to the texture quality and textual alignment. For each pair, three responses are collected, and the final result is determined by majority voting. For evaluation, we use $256$ random objects from the previously mentioned Objaverse subset.
As shown in Figure 5, our two-stage approach outperforms all the baselines. Figure 4 presents a qualitative comparison of our method with the three baseline approaches. The comparison against our single-stage approach confirms that the second stage is highly effective for further texture refinement.
Also, we provide the results of our method using the smaller version of DeepFloyd-IF that contains 400M parameters. We observe that our method still outperforms competing works.
Overall, our method achieves impressive quality, effectively producing detailed, seamless, and high-fidelity textures, as further confirmed by quantitative results.
5.3 Ablation study
We conducted a series of experiments to understand the role of the key design choices in our approach.
FID KID
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
CasTex (diffuse map only) 21.8 4.0
CasTex 20.5 3.1
Table 2: Ablation of Physically Based Rendering (PBR) components. Optimizing the full set of PBR components improves texture quality compared to synthesizing only the diffuse map.
Figure 6: Textures after the first and the second stage.
Do the PBR components improve the texture synthesis? PBR components enable the modeling of complex visual effects like reflections and material properties, avoiding baked-in lighting and improving realism.
We compare our full PBR-based approach with a baseline that generates only the diffuse map, one of the PBR components. As shown in Table 2, generating the full set of PBR maps leads to better FID and KID scores compared to generating the diffuse map alone. We provide visualization of PBR maps in Appendix E. Despite better metrics, a closer study of illustrations indicates that the method occasionally stuggles with baked-in lighting.
FID KID
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
Text2Tex 21.3 4.1
CasTex (Text2Tex init, $k_{d}$ frozen) 21.3 4.2
CasTex (Text2Tex init, $k_{d}$ trained) 20.5 3.9
Table 3: Ablation study of initializing our second stage with Text2Tex outputs to generate PBR textures. The results show that allowing the diffuse map ($k_{d}$) to be optimized during our second stage leads to significant improvements in both FID and KID metrics, while keeping the diffuse map frozen maintains the same FID but slightly increases KID. This suggests that our approach can effectively serve as an enhancement module for existing texture generation methods.
Can our second stage improve other methods? We hypothesize that our second stage in the PBR texture generation pipeline can improve the results of other texture synthesis methods. To validate this, we experiment with applying our second stage to well-initialized diffuse maps from another method.
As shown in Table 3, our second stage can effectively generate PBR textures when initialized with diffuse maps from Text2Tex. We explore two configurations: (1) where we keep the initial diffuse map fixed ($k_{d}$ frozen) and (2) where we fine-tune the diffuse map during PBR generation ($k_{d}$ trained).
The results show that when the diffuse map is kept frozen, the FID score remains unchanged with a slight increase in KID. This may be because freezing $k_{d}$ also preserves any baked lighting information present in the initial diffuse map, preventing our method from correcting for these lighting inconsistencies. In contrast, when we allow diffuse map optimization ($k_{d}$ trained), our approach achieves significant improvements in both metrics, yielding the best FID and KID. This demonstrates that enabling diffuse map refinement allows our method to correct lighting artifacts and better harmonize the entire PBR material set, highlighting the potential of our approach as a versatile enhancement module for various texture generation pipelines.
FID KID
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
Paint-it (Paint-it envlight) 25.2 6.7
Paint-it (proposed envlight) 24.5 6.4
CasTex (Paint-it envlight) 23.3 5.5
CasTex (proposed envlight) 20.5 3.1
Table 4: Ablation of environment lighting. We discovered that environmental lighting may play a crucial role in the plausibility of resulting physically-based textures. The default environmental lighting map in Paint-it presents an even gray lighting across all directions. The difference between a specular and a rough diffuse surface with such lighting is barely distinguishable. Instead, we propose an environmental lighting map with a single pronounced light source and augment lighting with horizontal rotations.
How does environment lighting affect results? Among recent works, Paint-it is the most similar to ours, as it also relies solely on SDS-based optimization. Unlike our approach, their rendering algorithm uses a brighter, uniform lighting environment map, whereas our setup combines a pronounced light source and a weaker ambient lighting component.
To understand the effect of environment lighting on SDS-based optimization, we evaluate both Paint-it’s lighting setup and our alternative configuration. As shown in Table 5, Paint-it achieves better results when using our lighting setup, slightly improving its FID and KID. Similarly, our methods performance degrades with Paint-it’s environment lighting. We hypothesize that, unlike uniform lighting, a single pronounce light source allows to highlight specular components of the PBR texture during training. The hypothesis and the above observations agree with the results from [13].
FID KID Runtime
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$) (seconds, $\downarrow$)
CasTex (DIP, XL) 23.2 4.6 229
CasTex (XL) 22.8 4.4 174
CasTex (DIP, XL+L) 21 3.8 862
CasTex (XL+L) 20.5 3.1 710
Table 5: Deep Image Prior ablation. Even though DIP plays a crucial role in texture synthesis with latent diffusion models, for the cascaded diffusion models the induced regularization is obsolete and slightly reduces the resulting scores. Also DIP-based variants are more time-consuming, while all experiments generate textures at 1024$\times$1024 resolution.
DIP versus explicit parameterization. Following the approach of our closest competitor, Paint-it, we integrate Deep Image Prior (DIP) [38] into our method and evaluate its impact on texture optimization (see Section 4.1 for details). As shown in Table 5, our DIP-based variant achieves results comparable to our standard explicit parameterization approach but performs slightly worse in both FID and KID metrics while being significantly slower. Since explicit parameterization already delivers strong results and works faster, the added complexity of DIP does not offer a clear advantage in our framework.
6 Conclusion
In this work, we proposed an optimization-based texture synthesis approach based on cascaded diffusion models. We motivated the design with an analysis indicating the limitations of latent diffusion models. Our method improves upon recent optimization-based texture synthesis approaches in terms of automated metrics and user study. Even though our model benefits from using physically-based textures, we still observed baked-in lighting and occasional inconsistencies between a texture and the corresponding model geometry. Future work may address the baked-in lighting artifacts of the aforementioned limitations of Score Distillation Sampling with latent models.
Acknowledgements
We would like to thank Roman Chulkevich for his assistance with the computing cluster and Ksenia Ushatskaya for her help with the visualization of the generated textures used in the paper teaser. Special thanks go to Maxim Kodryan, whose mere existence made this research journey considerably more enjoyable. This research was supported in part through computational resources of HPC facilities at HSE University.
Appendix A Evaluation Protocol Details
A.1 Blender parameters
We render images using Blender 3.3.21 with the Cycles rendering engine. Eval camera positions are taken from a sphere with a radius of $1.4$, where the spherical coordinates $\varphi$ (azimuth angle) and $\theta$ (polar angle) are derived from the setup described in Text2Tex. The HDRI light map from polyhaven used in our experiments is visualized in Figure 7(a); we set the environment map strength to $0.7$.
Figure 7: Used HDRI light maps: (a) for evaluation; (b) for our pipeline; (c) for Paint-it.
A.2 Details of User Study
Every assessor was asked to evaluate the setup shown in Figure 8.
Figure 8: Assessor’s evaluation setup. Professional assessors were tasked with evaluating the quality of the generated textures and selecting the texture that best aligned with the suggested prompt. To ensure unbiased evaluations, the order of the methods was randomly shuffled for each setup.
Appendix B Environment lights
For our texture generation setup, we utilize the HDRI light map visualized in Figure 7(b). In contrast, the Paint-it HDRI light map, shown in Figure 7(c), represents an alternative lighting configuration. All the HDRI light maps used in this work are publicly available for download from polyhaven.
Appendix C Texture Synthesis Results in Verying Setups
In this section, we provide textures obtained with different diffusion model combinations to illustrate the impact of model size and the role of super-resolution model.
Figure 9: Textures obtained with middle (M) and extra-large (XL) diffusion models along with their versions improved with large (L) super-resolution model on the second stage.
We also provide additional illustrations to compare our method to the baselines.
Figure 10: Additional samples from the Objaverse dataset.
Appendix D High frequency study in DIP
Our investigation into integrating Deep Image Prior (DIP) with our framework reveals certain limitations that affect its practical utility in texture synthesis. Despite being computationally more expensive, our experiments suggest that DIP may not provide sufficient benefits to justify its implementation costs.
As shown in Figure 11, spectral analysis conducted on the Objaverse dataset [12] demonstrates why the parameterization of DIP performs poorly compared to our standard explicit approach. The spectral plot (Figure 11(a)) reveals DIP’s significant deficiency in high-frequency components, visually translating to outputs with noticeably less textural detail and surface variation. To further illustrate this effect, we provide additional examples in Figure 11(b) using two cow prompts, which clearly demonstrate the degradation of the detail in the generation based on DIP.
This limitation in creating detailed textures further validates our choice of explicit parameterization, which achieves better visual quality and quantitative metrics with less computational cost.
Figure 11: Spectral and visual comparison of texture generation approaches. (a) Power spectrum analysis (log scale) computed on the Objaverse subset, showing CasTex + DIP’s reduced power in high-frequency components compared to standard CasTex . (b) Additional examples showing texture generation results for two test prompts. CasTex produces rich, detailed textures while CasTex + DIP generates simplified outputs lacking fine-grained details, directly corresponding to its high-frequency limitations observed in the spectral analysis.
Appendix E Visualization of PBR Texture Maps
Even though physically based texture improve the overall results, our method occasionally struggles to fully disentangle the lighting effects. We visualize a few examples in Figure 12. For instance, consider partially baked highlights on the coffee maker or the suitcase.
Figure 12: Visualization of the PBR textures. Lighter colors indicate higher metallic and roughness values for the corresponding texutre components.

### v2

@@
 1 Introduction
-The creation of 3D assets is a crucial part of modern computer graphics applications, with demand being supported by such industries as movie production, industrial design, and computer gaming, to name a few.
-High-quality assets should simultaneously meet aesthetic requirements and technical production requirements (i.e., high-quality mesh topology and limited polygon count).
-As a result, designing a 3D model is a labor-intensive creative task that involves multiple skills and is hard to automate due to the lack of high-quality data and occasionally difficulty in formalizing target criteria.
-At the same time, the rapid development of generative models in the visual domain has already tremendously impacted various creative domains and even extends to 3D graphics.
-The successes in 3D reconstruction [25, 1, 17] and foundation text-to-image model [33, 32] led fully-automated end-to-end 3d asset creation prototypes [29, 23, 20].
-Such solutions use the prior knowledge acquired by the diffusion model to generate 3D assets meeting aesthetic requirements.
+The creation of 3D assets is a central component of modern computer graphics, with demand driven by industries such as film production, industrial design, and computer gaming.
+High-quality assets should simultaneously meet aesthetic requirements and technical production requirements (e.g., clean mesh topology and limited polygon counts).
+As a result, 3D modeling remains a labor-intensive creative task that requires multiple expert skills and is difficult to automate due to the scarcity of high-quality data and the challenge of formalizing target criteria.
+Meanwhile, the rapid progress of generative models in the visual domain has already transformed many creative domains and is now extending into 3D graphics.
+Advances in 3D reconstruction [23, 1, 15] and text-to-image foundation models [31, 30] have enabled early propotypes for fully automated, end-to-end 3d asset generation [27, 21, 18].
+These systems leverage the prior knowledge acquired by the diffusion model to generate visually appealing 3D content.
 However, the resulting assets come in a volumetric representation and are often incompatible with modern graphics engines.
-Even with conversion tools for extracting meshes from volumetric representation, the assets end up having overly detailed geometry with occasional high-level geometric inconsistencies (e.g., multiple faces, as in the Janus effect).
-The aforementioned limitations may be critical when the technical requirements come into play.
-Texture synthesis is a more modest attempt to automate 3D asset creation, aiming to create a realistic texture for a given 3D shape.
-The task assumes that all of the technical requirements regarding the model geometry are already addressed by an artist.
-To meet modern industry standards, in our work, we specifically consider the synthesis of physically based textures [4].
-Such textures present a versatile instrument able to capture a diversity of materials and create realistically looking assets within varying lighting scenarios.
-Some early texture synthesis approaches involved the back-projection of synthetic images on 3D shapes.
-However, such approaches are incompatible with PBR textures as they only allow the capture of a fraction of material properties through RGB colors.
-As an alternative, we consider an optimization based approach that allows capturing the material properties with back-propagation through the rendering engine.
-In particular, we continue the line of work studying various applications of Score Distillation Sampling [29].
-We aimed to design a zero-shot texture-synthesis solution without relying on auxiliary 3D data and foundation model fine-tuning.
-However, we note that the texture synthesis setup unveils the limitations of Score Distillation Sampling applied in conjunction with modern diffusion models.
-Our observations go in line with the results of [41].
-We continue their analysis and attribute the effect to the specifics of latent diffusion models.
-As an alternative, we consider cascaded diffusion models that guide the images in RGB space rather than latent space.
-Our solution follows the pipeline of cascaded models.
-First, we generate a coarse texture using SDS guidance with a low-resolution model.
-Then we refine the texture using SDS guidance with a super-resolution model.
-Below, we summarize our main contributions.
+Even when volumetric outputs are converted into meshes, the geometry tends to be excessively detailed and may contain high-level structural errors (e.g., multiple faces as in the Janus effect).
+The limitations mentioned above may be critical when technical requirements must be met.
+Texture synthesis offers a more targeted and practical path toward automating parts of the 3D asset pipeline by generating realistic surface appearance for a given shape.
+In this work, we assume that the geometry is provided by an artist and focus on synthesizing physically based textures [4], which allow for expressive material modeling and realistic rendering under diverse lighting conditions.
+Early texture synthesis methods relied on back-projecting synthetic images onto 3D surfaces.
+However, such approaches are incompatible with PBR materials because RGB images capture only a limited subset of material properties through RGB colors.
+As an alternative, we consider an optimization-based framework that enables back-propagation through a differentiable renderer to recover full material parameters.
+In particular, we build on a line of recent work exploring Score Distillation Sampling (SDS) [27] for optimization-based generation.
+Our goal is to develop a zero-shot texture-synthesis solution that avoids reliance on auxiliary 3D data and foundation model fine-tuning.
+However, we observe that the texture synthesis setup unveils the limitations of Score Distillation Sampling applied with modern diffusion models.
+Consistent with [36], we find that SDS often drives optimization toward images that drift outside the natural image manifold.
+We expand upon this observation and attribute the effect to inherent properties of latent diffusion models.
+Motivated by this analysis, we explore cascaded diffusion models, which operate in RGB space rather than latent space.
+Our method follows the cascaded-model structure: we first generate a coarse texture using SDS guided by a low-resolution model, then refine it using SDS with a super-resolution model.
+In the following, we summarize our main contributions.
   • We study the limitations of Score Distillation Sampling when applied to latent diffusion models.
-Our analysis indicates that the optimal solutions of the algorithm include images outside of the natural image domain.
-  • Motivated by our analysis, we introduce a texture synthesis approach using Score Distillation Sampling with cascaded diffusion models.
-The proposed approach mimicks the diffusion model architecture and acts in two stages.
-The first stage generates a coarse texture, while the second stage improves overall quality of the texture obtained at the first stage.
-  • We show that our approach outperforms state-of-the-art optimization based texture synthesis solutions on the Objaverse dataset.
-Score Distillation Sampling in pixel space allows omitting additional regularizers required in previous works.
-Additionally, we report the scaling performance of our method with the diffusion model size, the impact of lighting conditions used during training, and the role of physically based textures.
+We show that the corresponding optimization procedure tends to converge to images outside the natural image domain.
+  • Based on these insights, we introduce CasTex, a texture-synthesis approach using Score Distillation Sampling with cascaded pixel-space diffusion models.
+The proposed method mimics the architecture of a diffusion model and acts in two stages.
+The first stage generates a coarse texture, while the second stage improves the overall quality of the texture obtained in the first stage.
+  • We demonstrate that CasTex outperforms state-of-the-art optimization-based texture synthesis solutions on the Objaverse dataset.
+Score Distillation Sampling in pixel space allows omitting additional regularizers required in previous works and enables efficient, high-resolution texture generation.
+Additionally, we report the scaling performance of our method with the diffusion model size, the impact of lighting conditions used during training, and the role of physically-based textures.
 2 Related Work
@@
 These methods often struggle with limited generalization and required extensive manual adjustments.
-Recent advances in large-scale text-to-image diffusion models [32, 33] transform texture synthesis, allowing more flexible and realistic results.
-However, different approaches vary in how they leverage diffusion models to incorporate 3D information and ensure global texture consistency.
+Recent advances in large-scale text-to-image diffusion models [30, 31] have transformed texture synthesis, allowing more flexible and realistic results.
+However, different approaches vary in how they leverage diffusion models to incorporate 3D information and ensure texture consistency.
 Diffusion-based texture generation.
@@
 SyncMVD [20] tackles this by generating each view independently while employing a synchronization module that aligns outputs in every denoising step of the diffusion model.
-Similarly, TexFusion [5] suggest a module that aggregates multi-view object predictions into a unified latent texture representation, iteratively refining it throughout the diffusion process to ensure consistency across the entire object.
-Another approach, inspired by MVDream [34], leverages consistent multi-view image generation as an initialization step for texture synthesis.
+Similarly, TexFusion [5] suggest a module that aggregates multiview object predictions into a unified latent texture representation, iteratively refining it throughout the diffusion process to ensure consistency across the object.
+Another approach, inspired by MVDream [32], leverages consistent multiview image generation as an initialization step for texture synthesis.
 Meta 3D TextureGen [2] follows this idea by generating four consistent views, which are projected onto a 3D surface to form an initial texture for subsequent refinement in UV space.
-MVPaint [8] extends this pipeline by introducing a spatial-aware 3D inpainting stage between projection and UV refinement, further enhancing coherence across different object views.
+MVPaint [8] extends this pipeline by introducing a spatially-aware 3D inpainting stage between projection and UV refinement, further enhancing coherence across different object views.
 Optimization-based methods.
 Optimization-based techniques refine a texture by optimizing an objective function that evaluates rendered views.
-Early approaches such as CLIP-Mesh [26] and Text2Mesh [24] use differentiable rendering and CLIP [30] as a guidance signal to iteratively update texture parameters.
-However, these methods often produce blurry textures due to the lack of high-frequency details.
-More recent works incorporate diffusion models to improve texture optimization.
-Latent-Paint [23], Fantasia3D [7], and Paint-it [41] combine differentiable rendering with the Score Distillation Sampling (SDS) technique [29], distilling knowledge from a diffusion model to refine textures.
-While this approach enhances realism, it often suffers from color oversaturation and high-frequency artifacts.
-FlashTex[13] mitigates some of these issues by employing a fine-tuned multi-view model to generate an initial textured view of the object, followed by an optimization stage using L2-loss and SDS for refinement.
+Early approaches such as CLIP-Mesh [24] and Text2Mesh [22] use differentiable rendering with CLIP [28] as a guidance signal to iteratively update texture parameters.
+While these methods are simple and training-free, they tend to produce blurred textures lacking high-frequency detail.
+More recent work incorporate diffusion models to improve texture optimization.
+Latent-Paint [21], Fantasia3D [7], and Paint-it [36] combine differentiable rendering with Score Distillation Sampling (SDS) technique [27], distilling knowledge from a diffusion model to refine textures.
+Although this approach enhances realism, it often suffers from color oversaturation and high-frequency artifacts.
+FlashTex[12] mitigates some of these issues by using a fine-tuned multiview model to generate an initial textured view of the object, followed by an optimization stage using L2-loss and SDS for refinement.
 In this work, we continue the line of work employing SDS.
-However, we attribute some of the artifacts to the underlying latent diffusion model and propose using a pixel-space model.
-Aiming at physically-plausible textures, we use a text-to-image model pretrained on natural images and resort from using synthetic data for multi-view model adaptation.
+However, we attribute some of the artifacts to the underlying latent diffusion model and propose using a pixel space model.
+Aiming at physically plausible textures, we use a text-to-image model pretrained on natural images and resort from using synthetic data for multiview model adaptation.
 PBR texture generation.
-Most of the works discussed above generate only RGB, yielding limited photo-realism.
-The works of Fantasia3D, Paint-it, and FlashTex propose to synthesize more advanced physically based textures [4], which are able to mimic different materials and look realistic under different lighting.
-We asses a similar approach to generation of PBR textures with a different generative model.
-3 Preliminaries
-In our work, we rely on a pretrained text-to-image model to generate textures.
-To apply a model designed for image generation to a different task, we use Score Distillation Sampling (SDS).
-Next, we describe the algorithm and the rendering algorithm required to apply it.
-3.1 Score Distillation Sampling for Gradient Based Generation
-SDS gained popularity in text-to-3D setup.
-In [29], the authors consider a differentiable rendering mechanism that produces a random view of a 3D model.
-In particular, they parameterize the model with a volumetric radiance field and render it using random lighting and camera position.
-SDS leverages a pretrained text-to-image diffusion model to guide the model parameters with gradient descent.
-Crucially, the mild assumptions on the rendering algorithm allow one to apply the method to other differentiable image generation mechanisms.
-In later works, Score-Distillation Sampling was successfully applied to other rendering functions and image parameterizations such as Gaussian Splatting [40, 9, 37], vector graphics [39, 14], and image rasterization with physically-based shading [20, 7, 41].
+Most of the works discussed above generate only RGB, limiting physical realism.
+Fantasia3D [7], Paint-it [36], and FlashTex [12] propose to synthesize more advanced physically based textures [4], which are able to mimic different materials and look realistic under different lighting.
+We pursue a similar goal but employ a different class of generative models and demonstrate the advantages of SDS in pixel space.
+Figure 1: Overall pipeline.
+Our method consists of two stages.
+In the first stage, given a 3D mesh and a text prompt, we use a differentiable rendering pipeline to generate random views of the model under varying lighting conditions.
+The texture, initialized randomly, is optimized using the SDS gradient $\nabla_{\theta}\mathcal{L}_{\text{SDS}}(\theta)$ (Eq.
+1).
+In the second stage, we refine the texture produced in stage one using SDS with a super-resolution diffusion model.
+For each camera view and lighting setup, we render two images: one using the fixed texture from the first stage, and one using the current texture being optimized.
+We denote the SDS gradient adapted to the super-resolution model as $\nabla_{\theta}\mathcal{L}_{SR-SDS}$.
+Using the former image as a conditioning input, we backpropagate the SR-SDS gradients through the latter image.
+3 Method
+We leverage a pretrained text-to-image diffusion model to synthesize high-quality textures.
+Because such models are trained for unconditional image generation rather than texture optimization, we employ Score Distillation Sampling (SDS) to obtain gradients compatible with our differentiable rendering pipeline.
+Specifically, we couple SDS with a cascaded diffusion model to refine textures across multiple resolutions.
+The overall architecture is illustrated in Figure 1.
+3.1 Preliminaries on Score Distillation Sampling for Gradient Based Generation
 SDS defines gradients used to update the parameter $\theta\in\Theta$.
-Assume a mapping $x_{0}=g(\theta,z)$ that takes the parameter and a random variable $c$ and produces an image $x_{0}$.
-Given a diffusion process with noisy images $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ at time $t$ and scale-parameter $\bar{\alpha}_{t}$, the approach leverages a pretrained diffusion model $\epsilon_{\phi}(x_{t};y,t)$ to define a pseudo-objective used to update $x_{0}$.
-[^1] Throughout the paper, we assume that the diffusion model $\epsilon_{\phi}$ incorporates the condition $y$ through classifier-free guidance.
-The pseudo-objective $\hat{x}_{0}^{t}=\tfrac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\phi}(x_{t},t,y)}{\sqrt{\bar{\alpha}_{t}}}$ estimates the appearance of the rendered image $x_{0}$ based on a noisy sample $x_{t}$ with a single step diffusion model prediction.
-The resulting gradient is estimated as follows:
+Let $x_{0}=g(\theta,c)$ denote the mapping from the parameter $\theta$ and a random variable $c$ to a rendered image $x_{0}$.
+Given a diffusion process in which noise images are generated as $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ at timestep $t$ with noise scale $\bar{\alpha}_{t}$, the method leverages a pretrained diffusion model $\epsilon_{\phi}(x_{t};y,t)$ to define a pseudo-objective that guides updates to $x_{0}$.
+[^1] Throughout the paper, we assume that the diffusion model $\epsilon_{\phi}$ incorporates the conditioning $y$ via classifier-free guidance.
+The pseudo-objective $\hat{x}_{0}^{t}=\tfrac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\phi}(x_{t},t,y)}{\sqrt{\bar{\alpha}_{t}}}$ provides a single-step estimate of the clean image $x_{0}$ from the noisy sample $x_{t}$ using the diffusion model’s prediction.
+The resulting gradient is estimated as
 $\nabla_{\theta}\mathcal{L}_{\text{SDS}}(\theta)=\mathbb{E}_{t,\epsilon,c}\left[w(t)(x_{0}-\hat{x}_{0}^{t})\frac{\partial x_{0}(\theta,c)}{\partial\theta}\right],$ (1)
-where $w(t)$ is an adjustable scalar weight parameter.
-Note that [29] formulate the method in noise prediction $\epsilon$-terms, while we introduce the equivalent formulation in $x_{0}$-terms.
-On the one hand, the gradient is essentially the mean-squared loss gradient $\mathbb{E}\left[w(t)\|x_{0}-\hat{x}_{0}^{t}\|^{2}\right]$ with the omitted dependence of $\hat{x}_{0}^{t}$ on $\theta$.
-On the other hand, the analysis in [29] demonstrates that the procedure minimizes $\mathcal{L}_{SDS}(\theta)=\mathbb{E}_{t,c}\operatorname{KL}\left(q(x_{t}\mid t)\|p_{\phi}(x_{t}\mid y,t)\right)$ using the score $\epsilon_{\phi}(x_{t};y,t)=\nabla_{x_{t}}\log p_{\phi}(x_{t}\mid y,t)$.
-Originally, SDS was applied to a pixel-space diffusion model that generates low-resolution images.
-However, pretrained pixel-space models are less common, and, as a result, applications of the method were primarily focused on latent diffusion models (LDMs).
-These models approximate the distribution of the low-dimensional embeddings of a high-resolution images rather than the images itself.
-Crucially, the mapping $\text{enc}(\cdot)$ that produces the embedding is differentiable, as a result, can be considered as a component of a function generating an image:
+where $w(t)$ is a scalar weight parameter.
+While [27] express SDS in terms of noise prediction $\epsilon$, we present an equivalent formulation in the $x_{0}$-domain.
+Viewed from one perspective, this gradient has the form of a mean-squared-error objective, $\mathbb{E}\left[w(t)\|x_{0}-\hat{x}_{0}^{t}\|^{2}\right]$, where the depencence of $\hat{x}_{0}^{t}$ on $\theta$ is ignored when computing the gradient.
+From another perspective, the analysis in [27] shows that SDS minimizes the divergence $\mathcal{L}_{SDS}(\theta)=\mathbb{E}_{t,c}\operatorname{KL}\left(q(x_{t}\mid t)\|p_{\phi}(x_{t}\mid y,t)\right)$ where the pretrained diffusion model provides the score $\epsilon_{\phi}(x_{t};y,t)=\nabla_{x_{t}}\log p_{\phi}(x_{t}\mid y,t)$.
+Originally, SDS was applied to pixel-space diffusion models that generate low-resolution images.
+However, pretrained pixel-space models are relatively rare, and consequently most applications of SDS have focused on latent diffusion models (LDMs).
+These models approximate the distribution of low-dimensional latent embeddings of high-resolution images rather than the images themselves.
+Importantly, the encoder mapping $\text{enc}(\cdot)$ that produces these embeddings is differentiable and can therefore be incorporated directly into the image-generation pipeline:
 $g_{\textit{latent}}(\theta,c)=\text{enc}(g_{\textit{pixel}}(\theta,c)).$ (2)
-The potential advantage of latent diffusion models is the ability to guide high-resolution images compared to pixel-space models.
-At the same time, it requires additional resources to run forward and backward passes of the encoder network.
-In Section 4, we analyze other implications of applying SDS in latent space rather than pixel space.
+A key advantage of LDMs is their ability to guide the synthesis of high-resolution images more efficiently than pixel-space models.
+However, this comes at the cost of additional computation due to the encoder’s forward and backward passes.
+In Section 4, we further examine the implications of applying SDS in latent space rather than pixel space.
+Figure 2: Score distillation sampling applied to a super-resolution diffusion model.
+The horizontal axis indicates optimization progress over time.
+Top row: the same image $g(\theta)$ is used both as a diffusion-model input $x$ and as the conditioning image $x_{cond}$ for super-resolution.
+This setup is unstable and causes the optimization to drift away from the original image.
+Bottom row: the same procedure, but with fixed super-resolution condition $x_{cond}=g(\theta_{0})$.
+The fixed condition anchors the optimization dynamics, preserving the structure of the initial image.
+3.2 Applying SDS with Super-Resolution Diffusion Models
+The main drawback of pixel-space diffusion models is that generating high-resolution images demands substantial computational resources.
+While latent diffusion models mitigate this cost by using a VAE [16] to encode images into a lower-dimensional latent space, pixel-space approaches typically rely on a cascaded generation process.
+In such pipelines, a diffusion model first produces a low-resolution image, and a subsequent super-resolution model upsamples it to the target resolution.
+In [27], the authors apply SDS only with a low-resolution diffusion model.
+However, because this model is trained on downsampled images, it often fails to recover fine-grained details and high-frequency texture.
+To achieve high-fidelity outputs, we adopt the architectural structure of cascaded diffusion models.
+Specifically, we first run SDS with a low-resolution diffusion model, and then refine the output using SDS in combination with a super-resolution model.
+In the text-to-image setting, the second-stage model takes as input a low-resolution image along with a text prompt and timestep.
+This setup naturally suggests two strategies for applying SDS during refinement.
+The first strategy is to use the rendered image $g(\theta,c)$ as the diffusion-model input in SDS, while providing a downsampled version of the same image as a conditioning signal.
+Intuitively, this approach iteratively enriches the texture by adding details that were missing in previous gradient steps.
+The second strategy is to freeze the parameters $\theta_{0}=\theta$ obtained from the first stage, then run SDS with a fresh copy of $\theta$ as the optimization target and treat $\theta_{0}$ as a fixed conditioning input.
+In this case, the first-stage result serves as an anchor that guides the gradient descent dynamics.
+In our experiments, the first strategy tends to be less stable and often drifts toward out-of-domain images, whereas the second strategy more reliably enhances resolution while preserving the details of the initial result.
+Figure 2 illustrates these observations.
 3.3 Differentiable Physically Based Shading
 Our texture synthesis approach applies SDS in a setting, where $g(\theta,z)$ renders an image of a fixed 3D model given a set of texture parameters $\theta$ and a random camera position and scene lighting specified by $z$.
-For rendering, we follow the pipeline proposed by [27].
-In particular, we take a differentiable rasterization algorithm based on the Nvdiffrast library by [19].
-For textures, we adopt the Disney-physically based material model from [4], allowing us to incorporate the obtained textures into popular graphics engines without further modifications.
-The texture $\theta=(\theta_{d},\theta_{r},\theta_{m},\theta_{n})$ consists of a diffuse component $\theta_{d}\in\mathbb{R}^{h\times w\times 3}$, a roughness map $\theta_{r}$ parameterizing a specular lobe and a metalness factor $\theta_{m}$ used to interpolate between plastic and metallic appearance.
-Finally, a normal map $\theta_{n}$ stores surface normal information, allowing to simulate additional geometric detail even within low-polygon models.
-To model lighting we incorporate cube environment map with a split sum approximation for the outgoing radiance integral approximation [16].
-Figure 1: Overall pipeline.
-Our method consists of two stages.
-On the first stage, given a 3D mesh and a prompt our method employs a differentiable rendering pipeline to generate random views of the model under various lighting.
-We update the randomly initialized model texture using Score Distillation Sampling (SDS).
-On the second stage, refine the texture from the first stage using SDS with a super-resolution diffusion model.
-For the same view and lighting, we render two frames: a frame with the fixed texture taken from the first step and a frame with a current texture.
-Using the former frame as the condition, we back-propagate the SDS gradients through the latter frame.
+Rendering follows the differentiable rasterization pipeline of [25] implemented using Nvdiffrast [17].
+Our textures follow the Disney physically based material model [4] with $\theta=(\theta_{d},\theta_{r},\theta_{m},\theta_{n})$ representing diffuse albedo, specular roughness, metalness, and a normal map, respectively.
+This parameterization ensures compatibility with standard real-time rendering engines.
+Lighting is provided via an environment map using a split-sum approximation of the outgoing radiance integral [14].
+4 Score Distillation Sampling Analysis
+In this section, we analyze Score Distillation Sampling (SDS) to motivate our proposed method.
+We argue that SDS applied to latent diffusion models is inherently prone to generating artifacts due to the ill-posed nature of the underlying optimization problem.
+Following common practice, we conduct the analysis in the simplified setting of image generation.
+We use a mapping $g_{id}(\theta,c)=\theta$ that produces a 2D image $\theta$ and while ignoring the stochastic variable.
+The goal is to recover images that are consistent with the diffusion model’s distribution using gradient descent.
+Intuitively, SDS gradients are expected to push generated images $g(\theta,c)$ toward the maniforld of natural images.
+Using the chain rule, these gradients also propagate to the parameter $\theta$.
+However, with latent diffusion models, guidance occurs in latent space: $g(\theta,c)=\text{enc}(g_{id}(\theta,c))$.
+As a result, the training signal must propagate through the encoder, placing additional constraints on the optimization trajectory.
+To isolate the behavior of SDS and exclude confounding factors such as model mis-specification or stochastic gradient noise, we consider a toy diffusion model that represents a degenerate distribution $\delta(x-x^{*})$ concentrated at a single image $x^{*}$.
+This setup removes issues arising from imperfect score estimation or excessive variance in Eq.
+1.
+For the degenerate distribution, the optimal score model reduces to a closed-form denoiser:
+$\epsilon_{degenerate}(x_{t})=\tfrac{x_{t}-\sqrt{\bar{\alpha}_{t}}x^{*}}{\sqrt{1-\bar{\alpha}_{t}}}.$ (3)
+With constant weights $w(t)=1$, the SDS gradient simplifies to the deterministic gradient of the mean squared error:
+$\nabla_{\theta}\mathcal{L}_{SDS}(\theta)=\nabla_{\theta}\|g(\theta)-x^{*}\|^{2}.$ (4)
+Thus, in this idealized setting we can write
+$\mathcal{L}_{SDS}(\theta)=\|g(\theta)-x^{*}\|^{2}$ (5)
+We optimize Eq.
+5 under various parameterizations and report the results in Figure 3.
+For pixel-space diffusion, where $g(\theta)=\theta$, gradient descent quickly converges to $\theta\approx x^{*}$ (top row, left).
+For latent-space model, we assume $x^{*}$ to be the latent code of the same image and minimize $\|\text{enc}(g(\theta))-x^{*}\|^{2}$.
+Although the gradient descent converges and the decoder maps $\text{enc}(\theta)$ to the original image (top row, right), the optimized parameter $\theta$ itself is a poor approximation to the original image (top row, middle).
+In other words, SDS with latent diffusion models frequently finds near-optimal latent solutions whose pre-images fall outside the natural image domain.
+This phenomenon does not significantly affect some text-to-3D tasks, where the optimization variable is not directly visualized.
+However, in the context of texture synthesis—where $\theta$ is the texture—these artifacts severely limit the applicability of SDS.
+To address this issue, [36] proposes parameterizing textures via a deep convolutional network, following the Deep Image Prior (DIP) framework [35].
+This parameterization acts as an implicit regularizer that discourages unnatural images.
+Although effective to some extent, it introduces an additional convolutional network into the pipeline and, as we show below, still leaves noticeable artifacts.
+The bottom row of Figure 3 illustrates SDS optimization with the DIP parameterization.
+Let the generated image be produced by a U-Net with parameters $\theta_{dip}$ and a fixed Gaussian noise input.
+We denote this differentiable generator as $g_{\text{UNet}}(\theta_{dip})$.
+We optimize Eq.
+5 under both pixel-space and latent-space models.
+The DIP parameterization substantially improves the solution for latent diffusion (middle row), reducing the artifacts seen with the explicit parameterization.
+However, even with DIP, SDS in latent space fails to match the reconstruction quality of directly optimizing the latent code (middle and right bottom row).
+We also experimented with additional regularization based on decoder outputs but observed no meaningful improvement over standard SDS.
+In contrast, pixel-space optimization achieves significantly better reconstruction in both explicit and implicit parameterizations.
 Figure 3: Optimization based image generation for a toy single-mode diffusion model.
-We study the effect of image parameterization and the difference between pixel and latent space diffusion models.
-Score Distillation Sampling (SDS) finds the distribution mode for both latent and pixel space diffusion models.
+We study the effect of image parameterization and the difference between pixel and latent-space diffusion models.
+Score Distillation Sampling (SDS) finds the distribution mode for both latent-and pixel-space diffusion models.
 Top row: the pixel space model achieves near-perfect reconstruction quality measured by PSNR.
@@
 Bottom row: The implicit regularization proposed in [36] significantly improves the results for SDS for the latent diffusion model, but still falls short of the image reconstruction given by the decoder or pixel space diffusion model.
-4 Methodology
-In the following section, we start with a motivating example and then describe the proposed method. The overall scheme is illustrated in Figure 1.
-4.1 Limitations of Score Distillation Sampling for Latent Diffusion Models
-In this section, for latent diffusion models, we argue that the procedure is prone to producing samples with artifacts due to the ill-posed nature of the task.
-Following the common practice, we convey the analysis on a task of image generation.
-Specifically, we use a mapping $g_{id}(\theta,c)=\theta$ that outputs a 2D image $\theta$ and does not rely on the second stochastic variable.
-The goal is to generate images on par with the diffusion model samples with a gradient descent algorithm.
-Intuitively, the SDS gradients guide images $g(\theta,c)$ towards the manifold of natural images.
-Using the chain rule, these gradients also propagate to the parameter $\theta$.
-However, in the case of latent diffusion models guidance happens within the latent space: $g(\theta,c)=\text{enc}(g_{id}(\theta,c))$.
-As a result, the training signal must propagate through the encoder.
-Next, to mitigate some of the potential issues of SDS in our analysis, we consider a toy diffusion model that represents a degenerate distribution $\delta(x-x^{*})$ concentrated at $x^{*}$.
-The problems may include a poorly trained score estimation model or the excessive variance of the stochastic gradient 1.
-For the degenerate distribution, we use the optimal model that separates the Gaussian noise from the original sample
-$\epsilon_{degenerate}(x_{t})=\tfrac{x_{t}-\sqrt{\bar{\alpha}_{t}}x^{*}}{\sqrt{1-\bar{\alpha}_{t}}}.$ (3)
-Note that for the model from Eq. 3 and constant weights $w(t)=1$ the gradient in Eq. 1 simplifies to the deterministic gradient of the mean squared error:
-$\nabla_{\theta}\mathcal{L}_{SDS}(\theta)=\nabla_{\theta}\|g(\theta)-x^{*}\|^{2}.$ (4)
-In other words, in the specified toy setup, we can assume
-$\mathcal{L}_{SDS}(\theta)=\|g(\theta)-x^{*}\|^{2}$ (5)
-To analyze SDS, we ran the optimization in various setups and report the results in Figure 2.
-For the pixel-space diffusion model with $g(\theta)=\theta$, gradient descent rapidly converges to $\theta\approx x^{*}$ (top row, left).
-For the latent-space model we assume $x^{*}$ to be the latent code of the same image and minimize $\|\text{enc}(g(\theta))-x^{*}\|^{2}$.
-Although gradient descent converges and the decoder maps $\text{enc}(\theta)$ to the original image (top row, right), the optimized parameter $\theta$ itself is a poor approximation to the original image (top row, middle).
-In practice, SDS with latent diffusion models finds globally optimal solutions that fall outside of the natural image domain.
-Notably, the above observation does not limit text-to-3D applications.
-However, in the above toy setup and texture synthesis domain the observed artifacts limit the applicability of SDS.
-To avoid the artifacts, a recent work [41] proposed to parameterize textures with a deep convolutional network.
-Their solution builds upon the observations of [38] and acts as an implicit regularizer.
-It helps avoiding solutions outside of the natural image domain, but requires running an additional convolutional and, as we show below, still leads to noticeable image artifacts.
-In the bottom row of Figure 2, we report SDS performance with an implicit image parameterization from [41].
-We denote the parameterization as DIP (Deep Image Prior), following [38].
-In this case, the image is an output of a UNet network with weights $\theta_{dip}$ and a fixed input sampled from the standard Gaussian distribution.
-In the context of the SDS, we can interpret the forward pass of the UNet as the differentiable rendering algorithm $g_{\text{UNet}}(\theta_{dip})$ constrained to generate natural images.
-We minimize Eq.
-5 for the implicit parameterization assuming both latent and pixel-space models.
-The implicit parameterization significantly improves the result for the latent diffusion model (middle row).
-The artifacts present in the explicit parameterization are less noticeable.
-Still, even with the implicit parameterization SDS in the latent space is unable to match the reconstruction quality of the reconstructed latent code (middle and right bottom row images).
-In our preliminary experiments, we tried to impose additional regularization on $\theta$ using the decoder outputs, but did not achieve noticeable improvements compared to the default SDS.
-Besides that, pixel-space optimization achieves significantly better result for both explicit and implicit image parameterization.
-Given the above observations, in our work, we use SDS with pixel-space models.
-4.2 Applying SDS with Super-Resolution Diffusion Models
-The downside of pixel-space diffusion models is that the model for high-resolution images require significant resources.
-While latent diffusion models employ a VAE [18] to encode a high-resolution image into a low-dimensional latent code, the typical solution for pixel-space models is to generate an image in a cascaded fashion.
-At the first stage, diffusion model generates a low-resolution image.
-During the second stage, a super-resolution model generates a high-resolution version of the image from the first stage.
-In [29], the authors run SDS only with a low-resolution model.
-However, the low-resolution model may be unable to recover fine-grained details and high-frequency features as its training domain was limited to down-scaled images.
-To achieve highly detailed textures, in our work we mimic the architecture of cascaded diffusion models.
-First, we run SDS with a low-resolution diffusion model.
-Then, we refine the output of the first stage with SDS combined with a super-resolution model.
-Figure 3: Score distillation sampling applied to a super-resolution diffusion model.
-Horizontal axis represents progress over time.
-Top row: the same image $g(\theta)$ used both as a diffusion model input $x$ and the condition $x_{cond}$ for super-resolution.
-The resulting procedure is unstable and significantly diverges from the original image.
-Bottom row: same procedure with fixed super-resolution condition $x_{cond}=g(\theta_{0})$.
-The condition anchors the dynamics to the initial image.
-In text-to-image setup, the second stage model takes a low-resolution image as an input along with a text prompt and time.
-This leads to the two natural ways of applying SDS to refine the first stage output.
-The first option is to pass the rendered image $g(\theta,c)$ as the diffusion model argument in SDS and a down-scaled copy of the image as a condition.
-Intuitively, in this case, the algorithm gradually updates the texture by adding the details lacking after the previous gradient step.
-The second option is to freeze the parameter $\theta_{0}=\theta$ obtained at the first stage and run SDS using a copy of $\theta$ as the new parameter and the previous $\theta_{0}$ as a condition.
-In this case, the first stage output in the condition acts as an anchor for the gradient descent dynamics.
-In our experience, the procedure in the first solution tends to be less stable and leads toward out-of-domain images, whereas the second approach tends to improve the resolution while preserving the details of the first stage result.
-We illustrate our observations in Figure 3.
-In summary, we propose a texture synthesis approach based on a cascaded modification of the SDS. Figure 1 is a schematic representation of the overall approach.
 5 Experiments
@@
 Also, we considered the setting with the smaller model IF(M) at the first stage to explore the effect of the diffusion model scale on the texture quality.
-We compared the proposed method with three baselines that represent various approaches to texture synthesis.
-First, Text2Tex [6] is a well-established back-projection-based method generating non-PBR textures.
-Second, Paint-it [41] is a simplistic optimization-based method using SDS.
+We compared the proposed method with four baselines that represent various approaches to texture synthesis.
+First, Text2Tex [6] is a well-established backprojection-based method generating non-PBR textures.
+Second, SyncMVD [20] is a recent method that enforces consistency by generating and synchronizing multiple views.
+Third, Paint-it [36] is a simplistic optimization-based method using SDS.
 The crucial distinction between our method and Paint-it is the underlying diffusion model and texture parameterization.
 Finally, FlashTex [12] is another recent approach using SDS.
-It addresses the limitations of previous methods by fine-tuning the model used for guidance on multi-view 3D data with controlled lighting.
+It addresses the limitations of previous methods by fine-tuning the model used for guidance on multiview 3D data with controlled lighting.
 Again, all of the baselines are based on latent diffusion models for texture synthesis.
@@
 Figure 4: Qualitative comparison.
-We synthesized textures for models from the Objaverse dataset using the proposed approach and a number of recent competing methods.
+We synthesized textures for models from the Objaverse dataset using the proposed approach and several recent competing methods.
 Our method generates seamless textures with softer colors compared with latent diffusion-based approaches.
 Figure 5: Human preference study comparing our method with competing optimization based and back-projection baselines.
-Size FID KID Runtime
-($\downarrow$) ($\downarrow$) (×$\times 10^{-3}$, $\downarrow$) (seconds, $\downarrow$)
-Text2Tex 2.5B 21.3 4.1 549
+Method Size FID KID Runtime
+($\downarrow$) ($\downarrow$) ($\times 10^{-3}$, $\downarrow$) (seconds, $\downarrow$)
+Non-SDS methods
+Text2Tex 2.5B 21.3 3.9 549
+SyncMVD 1.5B 20.3 2.7 55
+SDS methods
 Paint-It 1B 25.2 6.7 892
 FlashTex 1.3B 29.0 9.5 280
-CasTex (M) 400M 24.9 5.6 45
-CasTex (M+L) 1.6B 23.2 5.0 628
-CasTex (XL) 4.3B 22.8 4.4 174
-CasTex (XL+L) 5.5B 20.5 3.1 710
+CasTex (M) 400M 24.6 5.4 38
+CasTex (M+L) 1.6B 22.9 4.6 628
+CasTex (XL) 4.3B 22.4 4.2 174
+CasTex (XL+L) 5.5B 19.5 3.0 710
 Table 1: Comparison of FID and KID scores for different text-to-texture generation methods.
@@
 The two-stage cascaded diffusion approach achieved the best performance, outperforming all baselines.
-We illustrate the two stages in Figure 6 and provide additional illustrations in Appendix C.
-The second stage significantly refines textures, enhancing fine-grained details and coherence.
+We illustrate the two stages contrasting with the noisy results from latent diffusion and provide additional illustrations in Appendix C.
+The second stage significantly refines the textures, enhancing fine-grained details and coherence.
 Notably, the weak performance of FlashTex on the benchmark is the result of the limited generalization abilities of the Light ControlNet proposed by the authors.
 To our surprise, Text2Tex delivered competitive results.
-Even though the method is the oldest baseline, the back-projected photo-realistic images lead to photo-realistic textures and renders.
+Even though the method is the oldest baseline, the back-projected photo-realistic images lead to photo-realistic renders.
+SyncMVD also achieved strong scores, outperforming other baselines, which highlights the effectiveness of its multi-view synchronization approach.
 Our method also achieves competitive synthesis times, with the two-stage variant completing generation in 12 minutes per mesh – comparable to baselines while delivering better quality.
@@
 User preference study.
-In addition, we conduct a human study to compare the perceptual quality of our generated textures against baseline methods.
+In addition, we conduct a human study to compare the perceptual quality of our generated textures with the baseline methods.
 The evaluation is conducted by professional assessors, asked to make a decision between two rendered textures given a textual prompt.
 The decision is made according to the texture quality and textual alignment.
-For each pair, three responses are collected, and the final result is determined by majority voting.
+For each pair, three responses are collected and the final result is determined by majority vote.
 For evaluation, we use $256$ random objects from the previously mentioned Objaverse subset.
-As shown in Figure 5, our two-stage approach outperforms all the baselines.
-Figure 4 presents a qualitative comparison of our method with the three baseline approaches.
-The comparison against our single-stage approach confirms that the second stage is highly effective for further texture refinement.
-Also, we provide the results of our method using the smaller version of DeepFloyd-IF that contains 400M parameters. We observe that our method still outperforms competing works.
+As shown in Figure 5, our large two-stage approach outperforms all the baselines.
+Figure 4 presents a qualitative comparison of our method with the four baselines.
+The comparison against the first stage confirms that the second stage is highly effective for further texture refinement.
+In addition, we provide the results of our method using the smaller version of DeepFloyd-IF that contains 400M parameters. We observe that our method still outperforms most competing works.
 Overall, our method achieves impressive quality, effectively producing detailed, seamless, and high-fidelity textures, as further confirmed by quantitative results.
 5.3 Ablation study
-We conducted a series of experiments to understand the role of the key design choices in our approach.
-FID KID
-($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
-CasTex (diffuse map only) 21.8 4.0
-CasTex 20.5 3.1
+We conducted a series of experiments to understand the role of key design choices in our approach.
+Method FID KID
+($\downarrow$) ($\times 10^{-3}$, $\downarrow$)
+CasTex (diffuse map only) 21.2 3.7
+CasTex 19.5 2.4
 Table 2: Ablation of Physically Based Rendering (PBR) components. Optimizing the full set of PBR components improves texture quality compared to synthesizing only the diffuse map.
-Figure 6: Textures after the first and the second stage.
-Do the PBR components improve the texture synthesis?
-PBR components enable the modeling of complex visual effects like reflections and material properties, avoiding baked-in lighting and improving realism.
+Figure 6: Qualitative analysis of PBR components generated by CasTex.
+Lighter colors indicate higher metallic and roughness values for the corresponding texutre components.
+The microscope shows a clear separation between painted non‑metals and polished steel optics, with low roughness for the lenses and higher values for rubber and plastics.
+The banjo treats the wooden shell and drumhead as non‑metal while assigning metalness to the hardware and a semi‑gloss roughness to the lacquered wood.
+The binoculars further demonstrate semantic consistency by distinguishing matte rubber eyecups and strap from the painted housing and metallic fasteners.
+Across all three, the normal map contributes fine surface detail without altering the geometry.
+Do the PBR components improve textures?
+PBR components enable the modeling of complex visual effects like reflections and material properties, avoiding baked-in lighting and improving realism, as qualitatively demonstrated in Figure 6 (additional illustrations in Appendix E).
 We compare our full PBR-based approach with a baseline that generates only the diffuse map, one of the PBR components.
 As shown in Table 2, generating the full set of PBR maps leads to better FID and KID scores compared to generating the diffuse map alone.
-We provide visualization of PBR maps in Appendix E.
 Despite better metrics, a closer study of illustrations indicates that the method occasionally stuggles with baked-in lighting.
-FID KID
-($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
-Text2Tex 21.3 4.1
-CasTex (Text2Tex init, $k_{d}$ frozen) 21.3 4.2
-CasTex (Text2Tex init, $k_{d}$ trained) 20.5 3.9
+Method FID KID
+($\downarrow$) ($\times 10^{-3}$, $\downarrow$)
+Text2Tex 21.3 3.9
+CasTex (Text2Tex init, $k_{d}$ frozen) 20.9 3.7
+CasTex (Text2Tex init, $k_{d}$ trained) 19.7 3.6
 Table 3: Ablation study of initializing our second stage with Text2Tex outputs to generate PBR textures.
@@
 As shown in Table 3, our second stage can effectively generate PBR textures when initialized with diffuse maps from Text2Tex.
-We explore two configurations: (1) where we keep the initial diffuse map fixed ($k_{d}$ frozen) and (2) where we fine-tune the diffuse map during PBR generation ($k_{d}$ trained).
-The results show that when the diffuse map is kept frozen, the FID score remains unchanged with a slight increase in KID.
-This may be because freezing $k_{d}$ also preserves any baked lighting information present in the initial diffuse map, preventing our method from correcting for these lighting inconsistencies.
-In contrast, when we allow diffuse map optimization ($k_{d}$ trained), our approach achieves significant improvements in both metrics, yielding the best FID and KID.
-This demonstrates that enabling diffuse map refinement allows our method to correct lighting artifacts and better harmonize the entire PBR material set, highlighting the potential of our approach as a versatile enhancement module for various texture generation pipelines.
-FID KID
-($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
-Paint-it (Paint-it envlight) 25.2 6.7
-Paint-it (proposed envlight) 24.5 6.4
-CasTex (Paint-it envlight) 23.3 5.5
-CasTex (proposed envlight) 20.5 3.1
+We explore two configurations: (1) keeping the initial diffuse map fixed ($k_{d}$ frozen) and (2) fine-tuning the diffuse map during PBR generation ($k_{d}$ trained).
+The results show that even when the diffuse map is kept frozen, our approach improves both FID and KID scores over the original Text2Tex output.
+This improvement is likely limited because freezing $k_{d}$ also preserves any baked-in lighting information present in the initial diffuse map, preventing our method from correcting these inconsistencies.
+In contrast, when we allow diffuse map optimization ($k_{d}$ trained), our approach achieves a significant improvement in FID and a further, albeit smaller, improvement in KID.
+This shows that enabling diffuse map refinement helps to correct lighting artifacts and better harmonize the entire PBR material set, highlighting its potential as a versatile enhancement module for various texture generation pipelines.
+Studio light eval Paint-It light eval
+Method (Training envlight) FID KID FID KID
+($\downarrow$) ($\times 10^{-3}$, $\downarrow$) ($\downarrow$) ($\times 10^{-3}$, $\downarrow$)
+Paint-It (Paint-It) 25.2 6.7 26.8 6.2
+CasTex (Paint-It) 22.2 4.6 24.7 4.5
+Paint-It (Proposed) 24.5 6.4 26.9 6.6
+CasTex (Proposed) 19.5 2.4 24.1 4.4
 Table 4: Ablation of environment lighting.
-We discovered that environmental lighting may play a crucial role in the plausibility of resulting physically-based textures.
-The default environmental lighting map in Paint-it presents an even gray lighting across all directions.
-The difference between a specular and a rough diffuse surface with such lighting is barely distinguishable.
-Instead, we propose an environmental lighting map with a single pronounced light source and augment lighting with horizontal rotations.
+We investigate how different training and evaluation environment maps affect PBR texture quality.
+Both CasTex and Paint-It are trained on either the Paint-It or our Proposed maps.
+Performance is then evaluated under two distinct lighting conditions: our proposed Studio (which differs from both training maps) and the Paint-It maps.
+Our full pipeline (CasTex trained with the Proposed envlight) achieves the best results in both evaluation scenarios.
+Notably, CasTex also outperforms Paint-It even when both are trained and evaluated on the less expressive Paint-It map.
 How does environment lighting affect results?
 Among recent works, Paint-it is the most similar to ours, as it also relies solely on SDS-based optimization.
-Unlike our approach, their rendering algorithm uses a brighter, uniform lighting environment map, whereas our setup combines a pronounced light source and a weaker ambient lighting component.
-To understand the effect of environment lighting on SDS-based optimization, we evaluate both Paint-it’s lighting setup and our alternative configuration.
-As shown in Table 5, Paint-it achieves better results when using our lighting setup, slightly improving its FID and KID.
-Similarly, our methods performance degrades with Paint-it’s environment lighting.
-We hypothesize that, unlike uniform lighting, a single pronounce light source allows to highlight specular components of the PBR texture during training.
-The hypothesis and the above observations agree with the results from [13].
-FID KID Runtime
-($\downarrow$) (×$\times 10^{-3}$, $\downarrow$) (seconds, $\downarrow$)
-CasTex (DIP, XL) 23.2 4.6 229
-CasTex (XL) 22.8 4.4 174
-CasTex (DIP, XL+L) 21 3.8 862
-CasTex (XL+L) 20.5 3.1 710
-Table 5: Deep Image Prior ablation.
-Even though DIP plays a crucial role in texture synthesis with latent diffusion models, for the cascaded diffusion models the induced regularization is obsolete and slightly reduces the resulting scores.
-Also DIP-based variants are more time-consuming, while all experiments generate textures at 1024$\times$1024 resolution.
+Unlike our approach, their rendering algorithm uses a brighter, uniform lighting environment map for training, which can make it difficult to distinguish between specular and rough diffuse surfaces.
+Instead, we propose an environmental map with a single pronounced light source, a weaker ambient lighting component, and additional augmentations with horizontal rotations.
+For final evaluation, we also employ a distinct studio lighting environment, different from the one used for training (for more details, see Appendix A).
+To disentangle the effects of the model from the lighting strategy, we evaluated all combinations of training and evaluation setups against Paint-It in Table 4.
+Our full pipeline (CasTex with our proposed envlight) consistently achieves the best metrics across both evaluation scenarios, demonstrating its robust performance.
+Even when both models are trained and evaluated on the same, less expressive Paint-It environment map, CasTex outperforms Paint-It, highlighting an advantage of our model that stems from its design.
+Res. Method Runtime VRAM FID KID
+(min, $\downarrow$) (GB, $\downarrow$) ($\downarrow$) ($\times 10^{-3}$, $\downarrow$)
+1K CasTex (DIP) $\approx 14$ 20.7 20.8 3.7
+CasTex $\mathbf{\approx 12}$ 10.5 19.5 2.4
+2K CasTex (DIP) $\approx 24$ 52.3 21.1 3.9
+CasTex $\mathbf{\approx 12}$ 10.8 19.7 2.6
+4K CasTex (DIP) OOM OOM
+CasTex $\mathbf{\approx 12}$ 12 19.6 2.6
+[…]
 DIP versus explicit parameterization.
-Following the approach of our closest competitor, Paint-it, we integrate Deep Image Prior (DIP) [38] into our method and evaluate its impact on texture optimization (see Section 4.1 for details).
-As shown in Table 5, our DIP-based variant achieves results comparable to our standard explicit parameterization approach but performs slightly worse in both FID and KID metrics while being significantly slower.
-Since explicit parameterization already delivers strong results and works faster, the added complexity of DIP does not offer a clear advantage in our framework.
+Following the approach of our closest competitor, Paint-it, we integrate Deep Image Prior (DIP) [35] into our method and evaluate its impact on texture optimization (see Section 4 for details).
+We test our base model CasTex (XL+L) against its modification with implicit texture representation, CasTex (DIP, XL+L).
+Table 5 shows the results for quality, performance, and scalability across multiple resolutions.
+The analysis reveals two key findings.
+First, the DIP-based variant does not provide a significant quality benefit while being noticeably slower at all resolutions.
+Second, our method with explicit parameterization demonstrates remarkable scalability, maintaining a nearly constant runtime and only marginal VRAM growth.
+This confirms that for our cascaded approach, the added complexity of DIP is unnecessary for quality and scalability, making our explicit parameterization the best choice for practical applications.
 6 Conclusion
-In this work, we proposed an optimization-based texture synthesis approach based on cascaded diffusion models.
-We motivated the design with an analysis indicating the limitations of latent diffusion models.
-Our method improves upon recent optimization-based texture synthesis approaches in terms of automated metrics and user study.
-Even though our model benefits from using physically-based textures, we still observed baked-in lighting and occasional inconsistencies between a texture and the corresponding model geometry.
-Future work may address the baked-in lighting artifacts of the aforementioned limitations of Score Distillation Sampling with latent models.
+In this work, we introduced CasTex, an optimization-based texture synthesis method that leverages cascaded pixel-space diffusion models with explicit texture parameterization.
+We motivated our design through an analysis highlighting the limitations of latent diffusion models.
+Our method demonstrates improved performance over recent optimization-based texture synthesis approaches, as measured by automated metrics and user studies.
+While CasTex benefits from physically-based textures, we still observe baked-in lighting and occasional inconsistencies between textures and the underlying geometry.
+Future work may address baked-in lighting artifacts or the aforementioned SDS limitations in latent-space models.
 Acknowledgements
-We would like to thank Roman Chulkevich for his assistance with the computing cluster and Ksenia Ushatskaya for her help with the visualization of the generated textures used in the paper teaser.
+The paper was prepared within the framework of the HSE University Basic Research Program and was supported in part through computational resources of HPC facilities at HSE University.
+This research was also supported by Yandex Education.
+We would like to thank Roman Chulkevich and Viacheslav Kozyrev for their assistance with the computing cluster.
+We also thank Ksenia Ushatskaya for her help with the visualization of the generated textures used in the paper teaser.
 Special thanks go to Maxim Kodryan, whose mere existence made this research journey considerably more enjoyable.
-This research was supported in part through computational resources of HPC facilities at HSE University.
 Appendix A Evaluation Protocol Details
@@
 All the HDRI light maps used in this work are publicly available for download from polyhaven.
-Appendix C Texture Synthesis Results in Verying Setups
-In this section, we provide textures obtained with different diffusion model combinations to illustrate the impact of model size and the role of super-resolution model.
+Appendix C Texture Synthesis Results in Varying Setups
+In this section, we present textures obtained with different diffusion model combinations to illustrate the impact of model size and the role of the super-resolution module.
+The results are shown in Figure 9.
+We also provide additional examples comparing our method with the baselines in Figure 10.
+Finally, Figure 11 shows a textured model after the first and the second stage of using our cascaded pixel diffusion, compared to the noisy output of latent diffusion.
 Figure 9: Textures obtained with middle (M) and extra-large (XL) diffusion models along with their versions improved with large (L) super-resolution model on the second stage.
-We also provide additional illustrations to compare our method to the baselines.
 Figure 10: Additional samples from the Objaverse dataset.
+Figure 11: Textured model after the first and the second stage of our cascaded pixel diffusion, compared to the noisy result from latent diffusion.
 Appendix D High frequency study in DIP
@@
 For instance, consider partially baked highlights on the coffee maker or the suitcase.
+Figure 13 further evaluates prompt‑conditioned control on a fixed mesh: we change only the material description and generate the corresponding PBR maps.
+The generator consistently shifts the distributions of metalness, roughness, and normal amplitude in a physically meaningful way, indicating that it learns material semantics rather than merely recoloring the surface.
+Metals remain the most sensitive case, where map statistics can still correlate with lighting; we attribute this to limited light map diversity and leave stronger lighting disentanglement for future work.
+Figure 13: Prompt‑conditioned PBR textures demonstrating prompt‑level control over material properties.
+Top‑left (cork): high roughness variation, porous normal detail, near‑zero metalness.
+Top‑right (granite): stone‑like appearance from speckled albedo and fine normals; non‑metallic, mid–high roughness.
+Bottom‑left (soft‑touch plastic): uniform albedo, near‑zero metalness, satin roughness, low‑amplitude normals.
+Bottom‑right (steel): high metalness and lower roughness; residual noise and partially baked highlights remain—likely due to limited environment‑lighting variation; improving metal disentanglement is future work.
 Figure 14: Visualization of the PBR textures. Lighter colors indicate higher metallic and roughness values for the corresponding texutre components.